### v1

# Advancements in Constitutive Model Calibration: Leveraging the Power of Full-Field DIC Measurements and In-Situ Load Path Selection for Reliable Parameter Inference

Abstract
Accurate material characterization and model calibration are essential for computationally-supported high-consequence engineering decisions. Current characterization and calibration methods (1) use simplified test specimen geometries and global data, (2) cannot guarantee that sufficient characterization data is collected for a specific model of interest, (3) use deterministic methods that provide best-fit parameter values with no uncertainty quantification, and (4) are sequential, inflexible, and time-consuming.
This work brings together several recent advancements into an improved workflow called Interlaced Characterization and Calibration (ICC) that advances the state-of-the-art in constitutive model calibration. The ICC paradigm (1) employs tools to efficiently use full-field data to calibrate high-fidelity material models, (2) aligns the data needed with the data collected by adopting an optimal experimental design protocol, (3) quantifies parameter uncertainty through Bayesian inference, and (4) incorporates these advances into a quasi real-time feedback loop. The ICC framework is demonstrated here on the calibration of a material model using simulated full-field data for an aluminum cruciform specimen being deformed bi-axially. The cruciform is actively driven through the myopically optimal load path using Bayesian optimal experimental design, which selects load steps that yield the maximum expected information gain (EIG). To aid in numerical stability and preserve computational resources, the full-field data is dimensionally reduced via principal component analysis (PCA), and fast surrogate models which approximate the input-output relationships of the expensive finite element model are used. The tools developed and demonstrated here show that high-fidelity constitutive models can be efficiently and reliably calibrated with quantified uncertainty, thus supporting credible decision-making and potentially increasing the agility of solid mechanics modeling by enabling utilization of computational simulations at earlier stages of the design cycle.SAND2024-15320O
keywords:
Bayesian optimal experimental design , load path optimization , uncertainty quantification , constitutive model calibration , digital image correlation , principal component analysis , Gaussian Process surrogate modeling
journal: arXiv
[…]
[1]organization=Sandia National Laboratories, addressline=PO Box 5800, postcode=87185, city=Albuquerque, NM, country=USA
[…]
[figure]style=plain,subcapbesideposition=top,font=large
1 Introduction
Computational simulation is relied upon to understand complex engineering problems and provides essential information in decision-making. In finite element analysis of solid mechanics, constitutive relationships describe the response of a material to various external stimuli. Typically, these constitutive equations contain material-dependent parameters that are determined through material characterization and model calibration. Calibration and subsequent validation of the constitutive equations and parameters is an essential part of accurate computational simulations and structural analyses. Thus, there is a corresponding level of emphasis placed on the calibration process, both in terms of efficiency and accuracy, in order to establish confidence in predictions of the engineering problem at hand.
Traditional characterization and calibration procedures are often suboptimal in terms of efficiency and accuracy, which challenges the effective use of computational simulation for design purposes. The two processes (material characterization and model calibration) are typically performed sequentially and independent of each other, and it is not uncommon for multiple experimental campaigns to be needed in order to obtain the proper data for calibration [1]. In addition to the procedure, the type of data collected is an important consideration for calibration efficiency. Traditionally, tests which target specific stress states (i.e., uniaxial tension, notched tension, top hat shear, etc.) have been used for calibration. These tests provide limited information per specimen; therefore, large test matrices are needed which include various tests, sample geometries and cut directions in order to reveal complex material behaviors. As a result of the current characterization and calibration workflow in tandem with traditional testing procedures, the model calibration process is elongated and expensive, thus delaying the support of computational simulation for design and engineering decisions.
Several recent advancements have led to improvements in constitutive model calibration. For one, the use of digital image correlation (DIC) to collect rich full-field data through camera-based measurements has led to a paradigm shift in material characterization. Also known by the community as Material Testing 2.0 (MT2.0) [2], full-field DIC data collected from complex testing configurations (specimen shape plus loading) has greatly improved the information content that can be garnered from a single experiment for model calibration [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. Moreover, researchers have developed several different methods—including Chebychev and Zernicke polynomials [14, 15, 16, 17, 18, 19] and various types of wavelets [20, 21, 22]—for decomposing the field data into a spectral domain in order to reduce the dimensionality of the data while retaining its rich information content. However, selection of the proper experimental configuration, or design, which greatly influences the suitability of the data for model calibration, remains a challenging task. Thus, the second area of improvement for model calibration is the development of procedures for optimal experimental design within the context of solid mechanics. In general, there are two different approaches which have been adopted: 1) optimization of the geometry of the specimen [23, 24, 25, 10], or 2) optimization of the load path of deformation [26, 27]. Both approaches use various techniques to construct an experiment which will yield the most useful data for model calibration, and, consequently, improve constitutive model calibrations and predictions.
While the advancements made in material characterization and experimental design are significant, they have been, until this point, developed independently of each other. The contribution of this work is to drive forward the state-of-the-art in model calibration by bringing together these recent advancements into one improved workflow by 1) developing tools to efficiently use full-field data to calibrate high-fidelity material models and 2) aligning the data needed with the data collected by adopting the optimal experimental design protocol introduced in Ricciardi et al. [27]. The new workflow, which has been coined Interlaced Characterization and Calibration (ICC), actively controls the load path of a complex specimen in situ. The ICC framework was validated in [27] for a simplified exemplar problem of deforming a material point and calibrating with stress-strain data. This work extends the ICC framework from a material point to a complete structural problem of a cruciform specimen tested in a planar biaxial load frame. The ICC framework is demonstrated synthetically with the simulated deformation of the cruciform specimen in order to calibrate a plasticity model for an aluminum alloy. The development of the tools required for this demonstration are a critical stepping stone in moving towards utilization of the ICC framework with real-time data collection and model calibration.
The remainder of this paper is organized as follows: First, the components of the ICC framework as applied to load path selection of a cruciform specimen in a biaxial load frame are presented (Sec. 2). Next, a discussion of the experimental setup, constitutive model, statistical methods and dimension reduction of full-field data, as well as all other details of the framework are described in detail (Sec. 3). The ICC framework is demonstrated on an exemplar problem to calibrate an elasto-plastic model using synthetically generated full-field DIC and global load data (Sec. 4). Finally, various ICC algorithmic decisions and considerations for future work are discussed (Sec 5). In summary, this work brings together several recent advancements in model calibration into one enhanced calibration framework to improve both the efficiency and results of constitutive model calibration.
2 ICC Framework
In preparation for using the ICC framework with real-time data collection and model calibration, a simulated experiment of a cruciform specimen being deformed biaxially was used for tool development. Instead of simulating a traditional cruciform test, which applies equal loading on both arms and yields homogenous stress data in the central gauge region, the movement on each axis was isolated in order to introduce complex stress fields which are representative of real-world loading conditions. Thus, the experimental configuration in tandem with the complex geometry of the cruciform provides rich information on hardening and anisotropy.
The steps of the framework are illustrated in Fig. 1. Starting with plot (a), the experiment begins with an initial load step—deforming the cruciform specimen along one of the two axes while holding the other axis fixed (displacement = 0). DIC field data and global load data are collected in (b). The field data is then dimensionally reduced in (c) via principal component analysis (PCA) in order to make the algorithm more numerically stable. Furthermore, working in a reduced dimensional space aided in the construction of surrogate models which were pre-built. In (d), model calibration is performed via Bayesian inference, which yields a distribution that describes the parameter uncertainty. The expected information gain (EIG) is then optimized over the load step options (applying an increment of displacement along either of the two axes) to determine the next load step in (e). Finally, in order to efficiently perform the calibration with an expensive FE model, in (f), a surrogate replacement of the model is used (built off-line prior to entering the feedback loop) which is computationally cheap to evaluate. Additionally, in order to control the experiment, the infinite load path space is reduced to a finite load path tree—as seen in (f)—, which enables load step selection—details following. The feedback loop continues for some pre-determined number of cycles (equivalently, load steps) and yields a calibrated model with quantified uncertainty given the load path that is chosen. Each of the components which comprise the ICC framework in this setting are discussed in detail in Sec. 3.
Figure 1: The ICC framework is shown for a cruciform specimen being deformed in an actively controlled biaxial load frame. An initial displacement is imposed on the cruciform along a single axis while keeping the other axis fixed (a). Full-field DIC and global load data are collected (b). The high dimensionality of the DIC data is reduced via PCA (c). Given the collected data and any prior knowledge on the model parameters, Bayesian inference is used to update the knowledge about the parameters with quantified uncertainty (d). The next load step is chosen as the one which has the greater EIG (e). Prior to entering the feedback loop, a surrogate model is constructed to approximate the expensive FE simulation at the locations where data may be collected and organized into a load path tree (f).
In reality, the load path space from which the optimal load path is being selected is infinite when considering strain rate, step size, direction, etc. In order to enable the selection of an optimal load path in real-time, the decision space must be moved from infinite to finite dimensions. This is done by reducing the infinite-dimensional load path space of material deformation to a discrete space through the introduction of a load path tree.
The same graph structure that was introduced in [27] is adopted here. The graph structure is depicted in Fig. 2, which is a binary tree that stands in for the biaxial loading of the cruciform specimen. The BOED workflow begins at the root node with the cruciform in its undeformed state. Each parent node has two children nodes that correspond to applying a fixed displacement increment along either Axis A (the horizontal axis) or Axis B (the vertical axis) of the cruciform while holding the alternate axis fixed. The ICC framework is not restricted to this tree structure, but it was chosen for this demonstration as a simple starting point. Other more complex, and perhaps more appropriate, options are possible and are discussed Sec.5.
Figure 2: The load path tree reduces the infinite-dimensional load path space to finite dimensions. Here, a binary tree is shown which stands in for the deformation of a cruciform specimen in a biaxial load frame. The top node represents the specimen in its undeformed state. Each parent node has two children corresponding to applying an increment of displacement along either Axis A (horizontal) or Axis B (vertical) while holding the alternate axis fixed. In the example tree, four possible load paths (1-4) are shown, each consisting of two load steps.
In summary, the ICC framework utilizes a combination of complex specimen geometry, load path optimization, and full-field DIC calibration data, leading to fewer experiments and increased information content for model calibration compared to traditional approaches. As a result, well-calibrated constitutive models with quantified parameter uncertainty (a natural output of BOED) supports decision-making at earlier stages in the design cycle.
3 Methods
In this section, the various methods and techniques utilized to perform material characterization and model calibration in the ICC framework are discussed. The model of the physical system is discussed in Sec. 3.1, with details of the model form in Sec. 3.1.1, the cruciform and finite element analysis in Sec. 3.1.2 and the synthetic data generation in Sec. 3.1.3. A description of dimension reduction of the DIC data through PCA is found in Sec. 3.2, and details of the surrogate model construction are found in Sec. 3.3. Bayesian techniques for inference and experimental design are discussed in Sec. 3.4.1 and Sec. 3.4.2, respectively.
\@fb@secFB
3.1 Model of the Physical System
The selection of the proper model to represent the phenomena of interest is an important consideration and is crucial for producing credible predictions of material behavior. The process is simplified in this synthetic example as the same model being used to generate the data is also the one being calibrated. Thus, model-form error is removed from consideration, and in the absence of noise, the model fully captures the data generation process.
3.1.1 Material Model Form
For the current study, the calibration of an elastoplastic constitutive model is being explored. The (synthetic) experiments considered in this work are conducted at close to quasistatic rates and nominally room-temperature. As such, a rate and temperature-independent model is leveraged. Further, while non-proportional loadings are being probed, reverse and/or cyclic paths are not being pursued. Thus, plastic hardening is taken to be isotropic, and considerations of multiple and/or alternative hardening forms are left to future work.
The final consideration in the selection of the model is the yield function. In the current study, multiaxial loadings are being leveraged enabling the consideration of non-quadratic and/or anisotropic yield functions. Multiple yield functions in those classes have been used to probe the response of various aluminum alloys and corresponding forms [28, 29, 30, 31, 32]. Each model has a different degree of complexity and corresponding number of calibration parameters. As this effort represents the first attempt at the novel ICC calibration approach, the relatively simple Hosford yield function [33] is considered to enable testing of the proposed scheme. While isotropic, the Hosford surface needs only a single calibration coefficient (the exponent, $a$) with different strengths in shear and tension. The single parameter enables keeping the dimensionality of the calibration optimization problem appropriate while introducing a term sensitive to the yield surface shape. Further, the Hosford model has been shown to reasonably reproduce various responses in aluminum alloys (e.g.  [28, 30]).
Combining these various considerations, the constitutive response of interest is taken to be,
$\sigma_{ij}=\mathbb{C}_{ijkl}\left(\varepsilon_{ij}-\varepsilon_{ij}^{\text{p}}\right),$ (1)
in which $\sigma_{ij}$, $\mathbb{C}_{ijkl}$, $\varepsilon_{ij}$, and $\varepsilon_{ij}^{\text{p}}$ are the Cauchy stress, elastic stiffness (assumed isotropic), total strain, and plastic strain, respectively. The yield function, $g$, is then written,
$g=\phi(\sigma_{ij})-\bar{\sigma}(\kappa),$ (2)
with $\kappa$ being the isotropic hardening variable (found via the consistency condition), $\phi$ the effective stress describing the shape of the yield function, and $\bar{\sigma}$ the flow stress capturing the size of the yield surface. The effective stress is taken to be that of Hosford and given as,
$\phi(\sigma_{ij})={\frac{1}{2}[|\sigma_{1}-\sigma_{2}|^{a}+|\sigma_{2}-\sigma_{3}|^{a}+|\sigma_{3}-\sigma_{1}|^{a}]}^{\frac{1}{a}},$ (3)
where $\sigma_{1},~{}\sigma_{2}$ and $\sigma_{3}$ are the (ordered) principle stresses and $a$ is the exponent. As is commonly done with aluminum materials, a Voce-type saturation expression is used for the flow stress such that,
$\bar{\sigma}(\kappa)=\sigma_{y}+A\left(1-\exp(-n\kappa)\right).$ (4)
In Eqn. (4), $\sigma_{y}$, $A$, and $n$ are the initial yield stress, exponential modulus, and exponential coefficient to be found via calibration. An associative flow rule is also utilized such that,
$\dot{\varepsilon}_{ij}^{\text{p}}=\dot{\kappa}\frac{\partial\phi}{\partial\sigma_{ij}}.$ (5)
The described constitutive model has been implemented as part of the Library of Advanced Materials for Engineering (LAMÉ) [34] in the Sierra/SolidMechanics finite element code [35]. It is implemented via a traditional implicit, closest point projection return mapping algorithm [36, 37, 38].
The model parameters being inferred from the data are the initial yield stress $\sigma_{y}$ and isotropic hardening variables $A$ and $n$ from Eqn. (4) and the exponent $a$ from Eqn. (3). These parameters are collectively referred to as the unknown parameters such that $\boldsymbol{\theta}:=[\sigma_{y},\,A,\,n,\,a]^{T}$. The elastic properties – Young’s modulus, $E$, and Poisson ratio, $\nu$ – are taken to be known as they are well established.
3.1.2 Cruciform Specimen
The key design parameters for the cruciform geometry used in the ICC framework were (1) a multi-axial/ complex state of stress, (2) multiple load-steps before failure, and (3) no yielding in the arms (i.e. failure would occur in the central gauge area instead of the arms). Many cruciform geometries exist in the literature [39, 40, 41, 42, 43, 44, 45, 46, 47, 48], but these geometries are not always in strict alignment with the above criteria. The cruciform for this work was designed according to the criteria presented in Marmos et al. [47], which capitalizes on using re-entrant radii and a thinned gauge area. The used geometry is shown notionally in Fig. 3, with complete dimensions included in Fig. 1 in A. Further details and considerations of the specimen design can be found in [49].
With an assumed eight-fold symmetry, one-eighth of the cruciform geometry was modeled, and symmetry displacement boundary conditions were applied to preserve computational resources. These boundary conditions consisted of a constant, uniform displacement at the distal nodes of each arm along the coordinate direction for Axes A or B while the other degrees of freedom were traction-free. The mesh was comprised of hexahedral elements with linear shape functions and a selective-deviatoric integration rule. The elements in the center region were considerably smaller than those in the arms.
The deformation of the specimen was driven by displacement boundary conditions applied at a constant rate, and the displacement increment was 0.25 mm per load step applied to each arm. For example, load path 1 in Fig. 2 consists of two load steps where Axis A (the horizontal arms of the cruciform) is displaced by 0.5 mm (0.25 mm twice) while Axis B (the vertical arms of the cruciform) is held fixed in displacement control. Given the symmetry boundary conditions, the displacement of each load step equates to 0.5 mm total extension along a given axis.
[…]
[figure]style=plain,subcapbesideposition=top,font=small
[…]
Figure 3: Cruciform geometry and mesh. Left: Eight-fold symmetry was assumed such that only the opaque region was modeled. Displacement boundary conditions for Axis A (blue) and Axis B (red) were prescribed at the nodes at the end of each arm. The displacement surrogates predict values at nodes in the circular region (green). Right: The mesh was comprised of hexahedral elements of varying sizes.
An implicit, quasistatic finite-element analysis (FEA) was performed using Sierra/SM [35] with the given geometry and material model form (Sec. 3.1.1). An adaptive time-stepping algorithm was employed for the backward Euler time integrator used with the global equilibrium solution.The quantities of interest (QoIs) extracted from the FEA, which can in principal be computed from experimental measurements, include nodal displacements and resultant forces along each axis of the cruciform. The QoIs are returned for both the horizontal and vertical directional components and are denoted by X and Y, respectively. Thus the input-output relationship of the physical model is written as $g(\boldsymbol{\theta})=[\text{disp}_{i},\text{load}_{i}],i\in[X,Y]$. [^1] The horizontal and vertical directional components of displacement have traditionally been denoted by U and V, respectively; however, in order to avoid confusion with other notation in this paper, X and Y are used instead.
3.1.3 Synthetic Data Generation
To generate synthetic experimental data of the cruciform specimen, the FEA was run with the assumed-true set of model parameters, $\boldsymbol{\theta}^{true}$, reported in Table 1. The values in $\boldsymbol{\theta}^{true}$ were loosely based on a traditional calibration [50, 51] of the same material under consideration here, so that this synthetic demonstration would closely mimic the expected behavior of a physical experiment.
Parameter Symbol Value
[…]
Poisson’s ratio $\nu$ 0.33
Initial yield stress $\sigma_{y}$ 42.51 ksi (293.1 MPa)
Hardening modulus $A$ 13.63 ksi (94.0 MPa)
Hardening exponent $n$ 14.35
Hosford exponent $a$ 11.19
Table 1: Material model parameter values used for synthetic data generation
The number of load steps and the displacement increment per load step were chosen to keep the maximum total strain imposed on the cruciform within the plastic regime of deformation for an aluminum alloy, away from failure [50]. The synthetic experiment was carried out through 5 load steps with each load step imposing a displacement increment of 0.25 mm per arm. Thus a load path of 5 identical load steps, i.e. pulling on the same axis for the duration of the experiment, would result in a total displacement of 1.25 mm along that axis (and a displacement of 2.50 mm when considering symmetry) and 0 mm of displacement along the alternate axis.
To aid in the description of the simulated cruciform experiment, the displacements and resulting forces from two contrasting load paths are shown in Fig. 4. The left plot shows results from a load path pulling on Axis A for 5 consecutive load steps (AAAAA), and the right plot shows an alternating load path (ABABA). The resulting forces for each directional component are plotted vs. simulation pseudotime using parameter values from Table 1, and displacements are plotted on the right axis. The 5 diamonds correspond to nodes in the load path tree, which is where measurements were collected for calibration.
Two data types were extracted from the FEA to be used as the synthetic experimental data: these were 1) the nodal displacements (which simulate full-field DIC displacement measurements) collected from 1,022 nodes in the circular region of the cruciform geometry, indicated by the green nodes in Fig. 3, and 2) the resultant forces along each axis of the cruciform. The QoIs were stored at the end of each load step, which coincide with each node of the load path tree (Fig. 2, but extended to 5 load steps). Independent and identically distributed (i.i.d.) Gaussian noise was added to the synthetic data to mimic what is seen experimentally. For the forces, a noise variance of 582.11 $\text{N}^{2}$ was used and for the nodal displacements, a noise variance of 4 µm2 was used. In the remainder of this paper, ‘observed data’ refers to data that was synthetically generated.
An example of the synthetic displacement data for the AAAAA (Figs. 1 and 2) and ABABA (Figs. 1 and 2) load paths can be found in F and G, respectively.
[…]
[figure]style=plain,subcapbesideposition=top,font=small
[…]
Figure 4: The left plot shows the resulting global forces from following load path AAAAA and the right plot from following load path ABABA using parameter values from Table 1. Locations of the measured data points are shown with black diamonds, which correspond to the nodes in the load path tree. The right axes of the plots show the associated displacement for each case.
\@fb@secFB
3.2 Dimension Reduction of DIC Data
The simulated DIC measurements provide rich information content for calibration; however, the high-dimensional representation of DIC data poses a numerical challenge for calculating the EIG (Sec. 3.4.2) and likewise in the construction of the surrogate models built for the expensive finite element analysis (Sec. 3.3).
There are various methods to choose from when considering dimension reduction of full-field data. Methods based on Chebychev and Zernicke polynomials have been developed [14, 15, 16, 17, 18, 19]; however, these are often sensitive to missing data and difficult to perform on irregular geometries. There are also methods based on various types of wavelets [20, 21, 22], including the hierarchical polynomial wavelet decomposition (HPWD) method [52]. HPWD combines features from Alpert wavelets along with QR factorization to create an orthogonal basis for data representation. Thus, HPWD works directly on the domain, or point cloud, with no dependence on the field measurements (e.g. full-field DIC displacement data), and inherently, without consideration of the material model. Furthermore, HPWD does not need collocated points. HPWD can operate between different point clouds associated with the same geometry, for instance, nodes from an FE mesh from a simulation and a point cloud from experimental DIC data; it does not require data to be remapped from one point cloud to the other. Finally, there are singular value decomposition (SVD)-based techniques, such as principal component analysis (PCA) [53] or proper orthogonal decomposition (POD) [54], which require training data from the FEA simulation to obtain a transformation basis. HPWD and PCA were the two methods considered for this work as both handle missing data and irregular geometries well and are capable of capturing both large and small scale patterns in the data. After an analysis of the two methods for the specific case of dimension reduction for full-field data from a cruciform specimen, PCA was chosen as it yielded a smaller reconstruction error with fewer retained components (or modes). The remainder of this section will present the high-level procedure for this dimension reduction technique.
PCA is a method for dimension reduction which performs a linear transformation on a data set $A\in\mathbb{R}^{u\times v}$ in order to project it onto a new lower-dimensional orthogonal basis, a.k.a. principal components, such that the maximum amount of variance in the data is preserved. This dimension reduction is accomplished through a decomposition of the original matrix $A$ into three new matrices via SVD [55] such that $A=USV^{T}$. The maximum number of principal components that can be extracted from the data matrix is determined by the rank $r$ of matrix $A$; given a full-rank matrix, $r\coloneqq\text{rank}(A)=\min\{u,v\}$.
Fig. 5 shows a schematic representation of SVD for a rectangular matrix $A$, where $u<v$. The matrices $U\in\mathbb{R}^{u\times r}$ and $V^{T}\in\mathbb{R}^{r\times v}$ are orthonormal bases with columns and rows, respectively, consisting of the left and right singular vectors of $A$, and $S\in\mathbb{R}^{r\times r}$ is a diagonal matrix of singular values. The directions of the principal components are explained by the rows of $V^{T}$ and their magnitudes by the diagonal entries of $S$. The principal components are ordered according to the amount of variance they explain in the original data set from greatest to least. Thus, their relative importance is revealed, and a subset of components can be used to explain a given amount of variance in the original data set. By retaining a number $p<r$ of principal components, the data may represented in a reduced-dimensional space with very little information loss. Further details and nuances of PCA are left to [53].
Figure 5: Schematic representation of singular value decomposition (SVD) of a non-square matrix $A\in\mathbb{R}^{u\times v}$ with $u<v$. SVD factorizes the matrix into three new matrices $A=USV^{T}$, with $U\in\mathbb{R}^{u\times r}$, $S\in\mathbb{R}^{r\times r}$, and $V^{T}\in\mathbb{R}^{r\times v}$, where $r$ is the rank of $A$. The columns and rows of $U$ and $V^{T}$ contain the left and right singular vectors of $A$, respectively. The singular values are contained in $S$. By retaining a number of principal components $p<r$, matrix $A$ can be expressed in a reduced-dimensional space.
The schematic representation of SVD in Fig. 5 is reflective of the context in which PCA was exercised in this work. Training data for the PCA decomposition was generated by first obtaining samples of the unknown model parameters. For this task, Halton samples [56], which consist of a deterministic sequence of prime numbers to produce a space-filling design, were used. 400 training samples were generated within the bounds specified for the unknown parameters in Table 2, $\boldsymbol{\theta}^{hal}\in\mathbb{R}^{u\times D}$, where $D$ is the dimensionality of the unknown parameter vector $\boldsymbol{\theta}$ ($D=4$ in this case), and $u=400$—all other parameters were held fixed at the values recorded in Table 1. Each parameter sample was then used as input to the FE model, $g(\boldsymbol{\theta}^{hal})$, to produce corresponding output, which was the nodal displacements and resultant forces for both the $X$ and $Y$ directional components. However, PCA was performed only for the high-dimensional displacement field, not the resultant forces.
Parameter Symbol Bounds
[…]
Hardening modulus $A$ $[1.0,20.0]$ ksi
Hardening exponent $n$ $[0.5,20.0]$
Hosford exponent $a$ $[4.0,16.0]$
Table 2: Bounds for the unknown parameters.
The generated displacement data comprised matrix $A$, which was a data set of $u=400$ nodal displacement files which contained $v=1,022$ values each (green nodes in Fig. 3). The important first step was to make $A$ zero-mean by subtracting the mean of the data set such that, $A^{\prime}=A-A^{\text{mean}},\quad A^{\text{mean}}\in\mathbb{R}^{1\times v}$. Fig. 6 shows the PCA analysis for the $X$ directional component of displacement after a single load step applied along Axis A (which corresponds to Node 1 in Fig. 2). Note, separate PCA analyses were performed for each directional component of displacement ($X$ and $Y$) and for each node in the load path tree. A small portion of the training data set is shown in Fig. 6(a), and the retained principal components, $V^{*T}\subseteq V^{T}$, $V^{*T}\in\mathbb{R}^{p\times v}$, and singular values of $S$ from the SVD are shown in Fig. 6(b) and (c), respectively. The singular values decrease in magnitude quickly, indicating a waning level of information content in the corresponding principal components. A number of components $p=5$ was retained, thus, reducing dimensionality of the data set from $400\times 1,022$ to $400\times 5$ after performing the linear transformation $A^{\text{pca}}=A^{\prime}V^{*}\in\mathbb{R}^{u\times p}$. The retained principal components at Node 1 in the load path tree (Fig. 2) are shown for both $X$ and $Y$ in B.
Figure 6: PCA uses SVD to factor a data matrix into 3 new matrices $A=USV^{T}$ by which dimension-reduction proceeds; this approach was used to reduce the full-field displacement measurements from DIC (here, synthetically generate from FEA). The first step was to generate $u$ parameter samples, $\boldsymbol{\theta}^{hal}\in\mathbb{R}^{u\times D}$, and evaluate the FE model at the sampled parameters, $g(\boldsymbol{\theta}^{hal})$. The output was collected into the data matrix, $A$, in (a). The first five principal components in $V^{T}$ were retained to produce $V^{*T}$ and are shown in (b). The corresponding singular values of the retained components in $S$ are shown with red dots in (c), and the singular values of the components that were truncated are shown with blue dots. The images are shown for the decomposition of the X directional component after a single load step along Axis A (Node 1 in Fig. 2).
The original data set can be reconstructed from the retained components with $A^{\text{recon}}=A^{\text{pca}}V^{*T}+A^{\text{mean}}$. Fig. 7(a) shows an example reconstruction of one of the data sets which comprised $A$. The point-wise absolute difference between the reconstructed and original data file is shown in Fig. 7(b), which is very small relative to the magnitude of the measurements. Fig. 7(c) shows that with only 5 retained principal components, over $99.9\%$ of the information content in the original data set was preserved. Thus, retaining just the first few components successfully describes the majority of the variation in the data set and results in minimal information loss. Furthermore, there is a negligible benefit in retaining any additional components.
Figure 7: The quality of the reduced-dimensional representation of the displacement field through PCA can be visualized by (a) reconstructing the displacement field with the $p=5$ retained principal components, $A^{\text{recon}}$, and (b) calculating the point-wise absolute difference between the reconstructed field and real field, $|A^{\text{recon}}-A|$. The low error seen here is expected since the 5 principal components explain (or capture) over 99.9% of the variance in the original data set (c). Thus very little information was lost with the truncation of the remaining principal components. The example reconstruction is shown for the displacement field of the X directional component after a single load step along Axis A (Node 1 in Fig. 2). Note, the displacement control was prescribed at the end of the cruciform arms, thus resulting in a slightly lower maximum displacement in the region of interest.
The new orthogonal basis which is obtained can now be used to transform any new DIC file which is on the same grid as that used for training. In an experimental example, the DIC data would need to pre-processed before applying the PCA transformation in order to obtain measured values on an identical grid as that used for the PCA training. Since the same model is being used for the data generation process and model calibration in this case, no pre-processing of the data was necessary. Thus, as experimental data is collected (or simulated in this case), the PCA transformation can be used to express the high-dimensional data $A^{\text{new}}\in\mathbb{R}^{1\times v}$ in a low-dimensional space, $A^{\text{new}}V^{*}\in\mathbb{R}^{1\times p}$.
Dimension-reduction of high-dimensional data with minimal information loss is a crucial component of being able to use the highly-informative DIC data for model calibration. Numerical issues in calculating the EIG (Sec. 3.4.2) due to data dimensionality were resolved by representing the data in a low-dimensional space. Furthermore, the construction of the surrogate models, discussed subsequently in Sec. 3.3, was made viable with dimension reduction.
\@fb@secFB
3.3 Surrogate Model
The combined Bayesian calibration via the Laplace approximation (Sec. 3.4.1) and EIG calculation (Sec. 3.4.2) requires a minimum of $10^{4}$ model evaluations. [^2] This number may vary based on the EIG sampling strategy and other selected settings. Each run of the FEA conservatively takes 2 minutes to run, which equates to approximately 13 days to complete a single cycle of the ICC loop. Clearly, this computational demand is not feasible for the quasi-real-time setting the ICC framework is intended to operate within. Therefore, in order to cost-effectively perform calibration of the constitutive model within the ICC framework, the model evaluation must be reduced significantly, which is accomplished through the construction of a computationally inexpensive surrogate model replacement of the expensive FEA.
Fig. 8 shows the workflow for building a surrogate replacement for the FEA. The data needed to train and test the surrogate model was created by using a space-filling Halton sequence to generate parameter samples $\boldsymbol{\theta}^{hal}$ that served as input to the FEA—Fig. 8(a)—which then returned the quantities of interest (QoIs) as output—Fig. 8(b). The displacement values were then dimensionally reduced via PCA to obtain the corresponding singular values of the retained principal components—Fig. 8(c)—via the process described in Sec. 3.2. The resulting collection of input-output pairs constitute the training and test data for the surrogates—Fig. 8(d). The trained surrogate model then takes the model parameters $\boldsymbol{\theta}$ as input and returns an approximation to the displacement singular values and load; the displacement field may then be reconstructed from the singular values via $\text{disp}_{i}^{\text{recon}}=\text{disp}^{\text{PCA}}_{i}V^{*T}_{i}+\text{disp}^{\text{mean}}_{i}$.
Figure 8: A workflow showing how the surrogate models were built. The first step was to (a) generate $u$ parameter samples of $D$ dimensions ($D=4$) with a space-filling Halton sequence to serve as input to the FE model, $g(\boldsymbol{\theta}^{hal})$. The FE model yielded nodal displacements and load data for the horizontal and vertical directions, $i\in[X,Y]$ (b). The displacement data was reduced via PCA in (c) to yield a reduced-dimensional representation of displacement through singular values of the $p$ retained principal components. The input parameters, loads, and nodal displacement singular values (all marked by the green dashed rectangles) created the input-output combinations used to train the surrogate models in (d). A separate surrogate model was trained for each QoI, $h\in[\text{disp}^{\text{PCA}},\text{load}]$ and directional component $i\in[X,Y]$ at each node in the load path tree in Fig. 2.
Individual surrogate models were built for each QoI (displacement singular values and load values) and each directional component ($X$ and $Y$) for each node in the load path tree (besides the root node which represents the undeformed state) to approximately map constitutive model parameters to the QoIs, $\tilde{g}(\boldsymbol{\theta})\approx g(\boldsymbol{\theta})$. For the given load path tree, the total number of surrogates needed is $|\Xi|^{T}\times 2(p+1)$. In the adopted notation, $\Xi$ is the set containing the design options at each decision point (i.e. Ξ={Axis A,Axis $\Xi=\{\text{Axis A},\text{Axis B}\}$), and $|\Xi|$ is the cardinality of the set—equal to 2 in this example. $T$ is the number of load steps in the load path tree, and $p$ is the number of retained principal components. Multiplication by 2 is to account for both directional components (X and Y), and the addition of 1 accounts for the global load values. By building surrogates for each node—and inherently under the the prescribed boundary conditions described in Sec. 3.1.2 (i.e., 0.25 mm displacement increments)—, the history dependence which is inherent in elastoplastic models was accounted for, consistent with the approach taken in [27]. In reality, experimental boundary conditions may be more complex than those assumed here; this point is further discussed in Sec. 5. The surrogate model used in this work is a Gaussian process (GP) [57] with an anisotropic squared exponential kernel.
With the aid of a fast surrogate approximation of the expensive high-fidelity model, parameter inference and load step selection was able to be performed in a rapid manner. The cycle time for a single load step was reduced from nearly 2 weeks to just several minutes, thus significantly alleviating the ICC computational burden.
\@fb@secFB
3.4 Bayesian Techniques
A brief discussion of the Bayesian paradigm for inference and experimental design is contained in this section. The reader is referred to [27] for a complete discussion on the theory and tools found in this work.
3.4.1 Bayesian Inference
In each cycle of the ICC feedback loop, a single load step is applied to the specimen, and the collected data is used to perform model calibration. In lieu of a deterministic approach for model calibration, Bayesian inference is used, which provides an elegant way to quantify parameter uncertainty.
The first step is to model any knowledge regarding the unknown parameters of the forward model, $\boldsymbol{\theta}\in\boldsymbol{\Theta}\subset\mathbb{R}^{D}$, that is available a priori to observing data. In the adopted notation, the parameters $\boldsymbol{\theta}$ belongs to subspace $\boldsymbol{\Theta}$ of dimension $D$. Data $\boldsymbol{y}\in\boldsymbol{Y}\subset\mathbb{R}^{N_{obs}}$ is then collected which belongs to the data space $\boldsymbol{Y}$ with dimensionality $N_{obs}$, which is the number of measurements. The probability model for the data is denoted by $f\left(\boldsymbol{y}\>|\>\boldsymbol{\theta}\right)$, which describes the probability of observing data $\boldsymbol{y}$ given parameters $\boldsymbol{\theta}$, and for fixed $\boldsymbol{y}$, is regarded as the likelihood function. Once this data become available, the prior model is conditioned upon it and updated to the posterior distribution, $\pi(\boldsymbol{\theta}\>|\>\boldsymbol{y})$, via Bayes’ rule,
[…]
Thus, both the data and prior beliefs are taken into account when calculating the posterior distribution, which is, generally speaking, not available in closed-form due to the dimensionality of the parameter space. For this reason, simulation techniques such as Markov Chain Monte Carlo (MCMC) [58, 59, 60, 61, 62] are widely used to obtain an approximation of the posterior distribution through one or more of a variety of sampling algorithms. However, in practical applications for real-time decision-making, MCMC is prohibitively expensive. Thus, within the ICC framework, a Laplace approximation to the posterior distribution is adopted due to its relative speed and ease to compute.
Laplace’s approximation provides an analytical expression for the posterior distribution. Define $J(\boldsymbol{\theta})\coloneqq\log\pi(\boldsymbol{\theta}\>|\>\boldsymbol{y})$, and let $\hat{\boldsymbol{\theta}}$ be the parameter vector which maximizes $J(\boldsymbol{\theta})$, $\hat{\boldsymbol{\theta}}=\operatorname*{arg\,max}_{\boldsymbol{\theta}\in\boldsymbol{\Theta}}J(\boldsymbol{\theta})$—also known as the maximum a posteriori (MAP) estimate. The covariance of the approximated posterior is determined by the inverse of the Hessian, $\boldsymbol{H}$, of $J(\hat{\boldsymbol{\theta}})$, $\Sigma=\boldsymbol{H}_{J}(\hat{\boldsymbol{\theta}})^{-1}$. In this work, the Hessian is calculated through central finite difference. Thus, the Laplace approximation to the posterior is a Gaussian distribution characterized by the MAP parameter estimate and the local curvature at the MAP estimate, $\tilde{\pi}^{L}(\boldsymbol{\theta}\>|\>\boldsymbol{y})=\mathcal{N}_{D}(\hat{\boldsymbol{\theta}},\Sigma)$. This approach for approximating the posterior is very efficient and makes quasi-real-time step selection and model calibration a viable task.
3.4.2 Bayesian Optimal Experimental Design
The ICC framework actively drives the experiment through the load path that will yield informative calibration data by leveraging Bayesian optimal experimental design (BOED). Let $\boldsymbol{\xi}$ denote the load path of the specimen from start to finish and $\xi_{t}$ denote each load step from $t=1,\dots,T$, where $T$ is the total number of load steps in the load path ($T=5$ in this example). Also, let $\boldsymbol{y}_{t}$ be the data which is collected at load step $t$.
Within each cycle of the ICC framework, Bayesian inference is performed (Fig. 1d), which produces a posterior probability distribution that quantifies the parameter uncertainty conditioned on the data collected thus far. For example, after collecting data at load step $t$, the posterior is obtained via $\pi(\boldsymbol{\theta}\>|\>\boldsymbol{y}_{1:t},\boldsymbol{\xi}_{1:t})\propto f(\boldsymbol{y}_{1:t}\>|\>\boldsymbol{\theta},\boldsymbol{\xi}_{1:t})\pi(\boldsymbol{\theta})$. Next, load step $\xi_{t+1}$ must be selected (Fig. 1e). While this selection has traditionally been accomplished through subject matter expertise, BOED accomplishes this task in the ICC framework by taking into account previously collected data and the current state of knowledge about the unknown parameters to determine which load step, from a set of candidate load steps $\Xi$, will be the most informative.
Specifically, in this work, the decision is made through the calculation of the expected information gain (EIG), which is derived from the Kullback-Leibler (KL) divergence. The KL divergence provides a measure of the statistical distance between two probability distributions; in this setting, the two distributions of interest are the prior $\pi(\boldsymbol{\theta})$ and the posterior $\pi(\boldsymbol{\theta}\>|\>\boldsymbol{y},\xi)$ after performing candidate load step $\xi$. The candidate load steps are the branches in the load path tree, which here are applying either a displacement increment along Axis A (horizontal) or Axis B (vertical) of the cruciform specimen, $\xi\in\Xi=\{\text{Axis A},\text{Axis B}\}$. The optimal load step is then the one that has the greater expected KL divergence, which means it is expected to yield a posterior distribution that has the greater statistical distance from the prior, and therefore contains more information about the parameters. Since the KL divergence is calculated before the experiment is performed and the data collected, in practice, the expectation of the KL divergence over the marginal distribution of all possible outcomes is used to yield the EIG of candidate load step $\xi$,
$\mbox{EIG}(\xi)=\mathbb{E}_{\boldsymbol{y}\>|\>\xi}[D_{KL}(\pi\left(\boldsymbol{\theta}\>|\>\boldsymbol{y},\xi\right)\;\|\;\pi\left(\boldsymbol{\theta}\right))],$ (7)
where $D_{KL}(\cdot\;\|\;\cdot)$ denotes the KL divergence between the two given distributions.
After some manipulation, the final form of the EIG is,
[…]
where $f(\boldsymbol{y}\>|\>\xi)$ is called the evidence. [^3] The interested reader is referred to [27] for a more complete derivation of the EIG from the KL divergence. Since the EIG is intractable for high dimensions, a double-nested Monte Carlo estimator is used to approximate the EIG,
$\widehat{EIG}_{MC}(\xi)=\frac{1}{N}\sum_{i=1}^{N}\log\frac{f\left(\boldsymbol{y}_{i}\>|\>\boldsymbol{\theta}_{i,0},\xi\right)}{\frac{1}{M}\sum_{j=1}^{M}f\left(\boldsymbol{y}_{i}\>|\>\boldsymbol{\theta}_{i,j},\xi\right)},\quad\boldsymbol{\theta}_{i,j}\sim\pi(\boldsymbol{\theta}),\quad\boldsymbol{y}_{i}\sim f(\boldsymbol{y}\>|\>\boldsymbol{\theta}_{i,0},\xi).$ (9)
The approximation in Eqn. (9) uses samples from the likelihood and prior to evaluate the inner and outer sum of the Monte Carlo estimator. The quality of the estimate is largely determined by $N$ and $M$, the number of samples used to evaluate the outer and inner sums, which control the estimator variance and bias, respectively [63].
The next load step is then chosen as the candidate load step which maximizes the approximated EIG in Eqn. (9),
$\xi_{t+1}=\operatorname*{arg\,max}_{\xi}\widehat{EIG}_{MC}(\xi).$ (10)
In order to incorporate previously collected data into the EIG calculation, at each decision point, the prior in Eqn. (9) is defined as the posterior from the previous step. For example, when choosing load step $t+1$, $\pi(\boldsymbol{\theta})\coloneqq\pi(\boldsymbol{\theta}\>|\>\boldsymbol{y}_{1:t},\xi_{1:t})$, resulting in parameter samples being drawn from the calculated posterior distribution conditioned on data from load steps $1,\dots,t$: $\boldsymbol{\theta}_{i,j}\sim\pi(\boldsymbol{\theta}\>|\>\boldsymbol{y}_{1:t},\boldsymbol{\xi}_{1:t})$. To start the EIG process and determine the first load step from the undeformed state, either (a) the EIG may be calculated for the two candidate load steps by sampling from the prior, or (b) the first load step may be pre-determined, in which case there is no EIG calculation for the first load step. In this work, the second option was pursued. Finally, in this work, a myopic (a.k.a. greedy) BOED algorithm is used, meaning a locally optimal decision is made for each load step, as opposed to a globally optimal decision for the entire load path. This helps preserve computational resources as the EIG can be very costly to compute, and the cost would increase exponentially if calculating for the global optimum.
A challenge of computing the EIG is that numerical underflow may occur as a result of one or more factors: 1) if the prior distribution does not place a sufficient amount of mass on the posterior, leading to calculations for the evidence term (the inner loop of the EIG MC estimator) that are below machine precision, and 2) in the presence of high-dimensional data where the covariance-scaled squared difference of the measurements and model output—the exponential term in Eqn. (13)—is less than 1. The former was encountered and successfully handled through importance sampling (details in C). The latter case is observed if the displacement field is used for calibration without dimension reduction; however, this contributing factor was easily resolved with the PCA dimension reduction process discussed in Sec. 3.2. In combination, these two approaches were successful in resolving issues with underflow.
In this work, $M=1\times 10^{3}$ and $N=1\times 10^{4}$, and the initial load step was chosen to be $\xi_{1}=\text{Axis A}$. Pre-determining the first load step preserved computational resources as this reduced the number of surrogate models that needed to be built: they were only built for the portion of the tree which followed Node 1 in Fig. 2. Additionally, as was observed in [27], if there is no previously collected data (i.e. prior to the first load step), the EIG is determined solely by the prior modeling choice and physical model. So in cases such as this where the prior is only weakly informative, the EIG is unlikely to show strong preference between the two load steps. After the initial load step, the EIG was calculated to determine all subsequent load steps in the load path.
4 Synthetic Cruciform Exemplar
In this section, the exemplar problem is presented and discussed. The statistical model and other algorithmic settings within the ICC framework are discussed (Sec. 4.1), and results of the load path selection and calibration are presented (Sec. 4.2).
\@fb@secFB
4.1 Problem Setup
The exemplar problem explored in this work is the calibration of a constitutive model for a cruciform specimen made from an aluminum alloy being deformed in a biaxial load frame. The model form under consideration is the elastoplastic constitutive model described in Sec. 3.1 with unknown parameter vector $\boldsymbol{\theta}:=[\sigma_{y},\,A,\,n,\,a]^{T}$. These are the initial yield stress $\sigma_{y}$ and the isotropic hardening variables $A$ and $n$ from Eqn. (4) and the exponent $a$ from Eqn. (3). This section is used to describe the statistical modeling decisions and other algorithm settings specific to this exemplar problem.
The parameters were constrained to take on values within the bounds specified in Table 2, which are the same as those use for surrogate training. The prior model was chosen to be a truncated normal probability distribution (TN) which enforced these bounds during the calibration,
$\pi(\boldsymbol{\theta}_{d})=TN\left(\boldsymbol{\theta}_{d};\mu_{(\boldsymbol{\theta}_{d})},\delta^{2}_{(\boldsymbol{\theta}_{d})},lb_{(\boldsymbol{\theta}_{d})},ub_{(\boldsymbol{\theta}_{d})}\right),\quad d=1,\ldots,D.$ (11)
The distribution means $\mu$ were chosen to yield an expected value within the bounds, and the variances $\delta^{2}$ were chosen to be moderately diffuse. The support of the distribution, defined by $lb$ (the lower bound) and $ub$ (the upper bound), was chosen to be the same as that used for the surrogate training. All distribution parameters are detailed in Table 3.
Parameter $\boldsymbol{\mu_{\theta}}$ $\boldsymbol{\delta^{2}_{\theta}}$ $\boldsymbol{lb}$ $\boldsymbol{ub}$
[…]
$A$ (ksi) 10.0 225.0 1.0 20.0
$n$ 10.0 225.0 0.5 20.0
$a$ 10.0 225.0 4.0 16.0
Table 3: Means $\mu_{\theta}$, variances $\delta^{2}_{\theta}$ and bounds $lb$ and $ub$ for the truncated normal prior model Eqn. (11) used in the exemplar problem. Values for parameters $\sigma_{y}$ and $A$ have units of ksi.
Simulated experimental data was generated by running the expensive FEA under assumed true parameter values $\boldsymbol{\theta}^{true}$ (detailed in Table 1) with added Gaussian error. This synthetic data is used here for model calibration to demonstrate the ICC framework under controlled conditions in preparation for later using it for real, physical experiments. The data consisted of the nodal displacements, dimensionally reduced via PCA to yield the singular values, and global load values for both the $X$ and $Y$ components, $\boldsymbol{y}=[\text{disp}^{\text{PCA}}_{i},\text{load}_{i}],\,i\in[X,Y]$, where $\text{disp}^{\text{PCA}}_{i}=\text{disp}_{i}V^{*}_{i}$ and $V^{*}_{i}$ is the matrix containing the retained principal components for directional component $i$. The FEA model was evaluated and measurements collected at the end of each load step, corresponding to the nodes in the load path tree.
The data generation process was modeled under the assumption that the observed data is equal to the surrogate model output (which consists of the dimensionally reduced nodal displacements and load values) with added Gaussian error $e$,
$\boldsymbol{y}_{h,i}=\tilde{g}(\boldsymbol{\theta})_{h,i}+e_{h,i},\qquad e_{h,i}\sim\mathcal{N}_{N_{h,i}}\left(0,\Psi_{h,i}\right),\,h\in[\text{disp}^{\text{PCA}},\text{load}],\,i\in[X,Y],$ (12)
where $\boldsymbol{y}_{h,i}$ and $\tilde{g}(\boldsymbol{\theta})_{h,i}$ are the data the surrogate replacement of the forward model for each QoI and directional component evaluated at $\boldsymbol{\theta}$, and $N_{h,i}$ is the number of measured observations for each QoI and directional component: $N_{\text{disp}^{\text{PCA}}}=5$ and $N_{\text{load}}=1$—the same for both $X$ and $Y$. The equivalent likelihood function for the expression in Eqn. (12) is a multivariate normal distribution centered at the surrogate replacement of the forward model with a covariance matrix $\Psi_{h,i}$ of size $N_{h,i}\times N_{h,i}$,
$f\left(\boldsymbol{y}_{h,i}\>|\>\boldsymbol{\theta},\boldsymbol{\xi}\right)$ $=\mathcal{N}_{N_{h,i}}\left(\boldsymbol{y}_{h,i};\tilde{g}(\boldsymbol{\theta})_{h,i},\Psi_{h,i}\right),$ (13)
$=(2\pi)^{-N_{h,i}/2}\det\left(\Psi_{h,i}\right)^{-1/2}\exp\left(-\frac{1}{2}\left(\boldsymbol{y}_{h,i}-\tilde{g}(\boldsymbol{\theta})_{h,i}\right)^{T}\Psi_{h,i}\left(\boldsymbol{y}_{h,i}-\tilde{g}(\boldsymbol{\theta})_{h,i}\right)\right),$
$h\in[\text{disp}^{\text{PCA}},\text{load}],\,i\in[X,Y].$
The observation noise variance is assumed to be independent and identically distributed (i.i.d.) such that $\Psi_{h,i}=\psi^{2}_{h,i}\mathbb{I}_{N_{h,i}}$, where $\mathbb{I}_{N_{h,i}}$ is an $N_{h,i}\times N_{h,i}$ identity matrix and $\psi^{2}_{h,i}$ is the observation noise variance—assumed to be known and fixed at $\psi^{2}_{\text{disp}}=4\mu\text{m}^{2}$ and $\psi^{2}_{\text{load}}=582.11\,\mbox{N}^{2}$ for both the $X$ and $Y$ directional components. Note, the experimental noise variance for the displacement was defined in the field space and then transformed to account for the linear transformation of the displacement field via PCA, $\Psi_{\text{disp}^{\text{PCA}},i}=V^{*T}_{i}\Psi_{\text{disp},i}V^{*}_{i},\,i\in[X,Y]$—additional details may be found in D. The noise level was chosen to be on the same order as that expected in experimental data. The likelihood is written as a product over the 4 data sets contained in $\boldsymbol{y}$ (load and displacement for the $X$ and $Y$ directional components) under the assumption of independence,
$f\left(\boldsymbol{y}\>|\>\boldsymbol{\theta},\boldsymbol{\xi}\right)=\prod_{h}\prod_{i}f\left(\boldsymbol{y}_{h,i}\>|\>\boldsymbol{\theta},\boldsymbol{\xi}\right),\quad h\in[\text{disp}^{\text{PCA}},\text{load}],\quad i\in[X,Y].$ (14)
The ICC algorithm was carried out to a total of $T=5$ load steps, with each load step applying a displacement increment of 0.25 mm along the specified axis. The first load step was pre-determined to be $\xi_{1}=\text{Axis A}$. After the initial load step, the EIG was calculated to determine all subsequent load steps in the load path—$[\xi_{2},\xi_{3},\xi_{4},\xi_{5}]$.
\@fb@secFB
4.2 Results
Model calibration was first performed with all possible load paths with an initial load step along Axis A for a total of 16 calibrations with pre-determined load paths. For ease of reading, A is used to denote a load step along Axis A (horizontal direction) and B for Axis B (vertical direction). Table 4 reports the expected values, $\mathbb{E}_{\boldsymbol{\theta}\>|\>\boldsymbol{y},\boldsymbol{\xi}}$, marginal variances, $\mathbb{V}_{\boldsymbol{\theta}\>|\>\boldsymbol{y},\boldsymbol{\xi}}$ and 95% credible intervals (CI) for parameter vector $\boldsymbol{\theta}=[\sigma_{y},A,n,a]$ for all 16 calibrations, with $\sigma_{y}$ and $A$ being reported in units of ksi. The marginal variance is the variance (or uncertainty) of one variable when looked at in isolation, without consideration of the others, and the 95% CI is the interval in which the parameter has a 95% probability to fall within. Since a Laplace approximation to the posterior distribution was used, the marginal posterior distributions are symmetric, and the 95% CIs are reported with a single value, $\zeta$, where the expected value plus and minus $\zeta$ defines the interval: $95\%\,\text{CI}=\mathbb{E}_{\boldsymbol{\theta}\>|\>\boldsymbol{y},\boldsymbol{\xi}}\,\pm\zeta$. Additionally, the collective parameter uncertainty, reported as the determinant of the posterior covariance $\det(\Sigma)$ (a.k.a. the generalized variance), is included.
The resultant parameter expected values were very similar for each load path. Specifically, $\mathbb{E}_{\sigma_{y}\>|\>\boldsymbol{y},\boldsymbol{\xi}}$ for the initial yield stress was in agreement to the tenth for all cases, and the expected value of all other parameters were in close agreement to the tenth. The marginal variances for $\sigma_{y}$ and $a$ were similar for all cases; however, there were orders of magnitude difference in the marginal variances reported for $A$ and $n$ among the different cases (although still on a relatively small scale). There was likewise orders of magnitude difference in the reported generalized variance among the different load paths. A similar trend was seen in the 95% CI; the interval was similar for parameters $\sigma_{y}$ and $a$ for all cases, but showed more deviation case to case for parameters $A$ and $n$. Thus parameters $A$ and $n$ were the most sensitive to the applied load path and resulting calibration data. In summary, the pre-determined AAAAA load path exhibited the least amount of parameter uncertainty, expressed through the various reported uncertainty metrics, thus demonstrating the highest confidence in the parameter values; likewise, the ABABA load path had the greatest parameter uncertainty.
Load path $\boldsymbol{\mathbb{E}_{\boldsymbol{\theta}\>|\>\boldsymbol{y},\boldsymbol{\xi}}}$ $\boldsymbol{\mathbb{V}_{\boldsymbol{\theta}\>|\>\boldsymbol{y},\boldsymbol{\xi}}}$ $\zeta$ Gen. Var.
$\boldsymbol{(1\times 10^{-4})}$ $\boldsymbol{(1\times 10^{-14})}$
[…]
AAAAB [42.5, 13.6, 14.4, 11.1] [ 2.0, 17.9, 84.7, 4.9 ] [0.03, 0.08, 0.18, 0.04] 1.8
AAABA [42.5, 13.6, 14.3, 11.2] [ 1.9, 24.5, 112.5, 5.2 ] [0.03, 0.10, 0.21, 0.04] 3.7
AAABB [42.5, 13.7, 14.2, 11.1] [ 1.7, 50.2, 143.2, 6.2 ] [0.03, 0.14, 0.23, 0.05] 10.0
AABAA [42.5, 13.5, 14.6, 11.2] [ 1.8, 15.1, 87.3, 6.3 ] [0.03, 0.08, 0.18, 0.05] 4.5
AABAB [42.5, 13.5, 14.5, 11.2] [ 1.8, 94.8, 299.6, 6.0 ] [0.03, 0.19, 0.34, 0.05] 35.1
AABBA [42.5, 13.6, 14.4, 11.1] [ 3.0, 63.8, 238.6, 7.9 ] [0.03, 0.16, 0.30, 0.06] 51.9
AABBB [42.5, 13.7, 14.3, 11.2] [ 2.5, 14.4, 96.4, 7.3 ] [0.03, 0.07, 0.19, 0.05] 6.8
ABAAA [42.5, 13.6, 14.4, 11.3] [ 1.4, 11.7, 69.3, 11.2 ] [0.02, 0.07, 0.16, 0.07] 5.4
ABAAB [42.5, 13.3, 14.9, 11.2] [ 2.3, 114.4, 484.6, 8.4 ] [0.03, 0.21, 0.43, 0.06] 55.5
[…]
ABABB [42.5, 13.5, 14.7, 11.2] [ 1.9, 26.6, 131.0, 13.0 ] [0.03, 0.10, 0.22, 0.07] 48.3
ABBAA [42.5, 13.6, 14.5, 11.2] [ 1.4, 31.7, 96.9, 10.8 ] [0.02, 0.11, 0.19, 0.06] 29.0
ABBAB [42.5, 13.8, 14.2, 11.2] [ 1.7, 168.6, 442.2, 7.1 ] [0.03, 0.25, 0.41, 0.05] 92.9
ABBBA [42.5, 13.6, 14.4, 11.2] [ 1.9, 108.6, 379.1, 7.5 ] [0.03, 0.20, 0.38, 0.05] 29.0
ABBBB [42.5, 13.6, 14.3, 11.2] [ 1.3, 8.1, 47.6, 9.9 ] [0.02, 0.06, 0.14, 0.06] 2.49
Table 4: Posterior summaries for 16 pre-determined load paths, each starting within an initial A load step. The parameter expected values, $\mathbb{E}_{\boldsymbol{\theta}\>|\>\boldsymbol{y},\boldsymbol{\xi}}$, marginal variances, $\mathbb{V}_{\boldsymbol{\theta}\>|\>\boldsymbol{y},\boldsymbol{\xi}}$, and 95% CI are reported for parameter vector $\boldsymbol{\theta}=[\sigma_{y},A,n,a]$ ($\sigma_{y}$ and $A$ reported in units of ksi) along with the generalized variance, which is the determinant of the posterior covariance. The 95% CI is represented as $\mathbb{E}_{\boldsymbol{\theta}\>|\>\boldsymbol{y},\boldsymbol{\xi}}\,\pm\zeta$, where $\zeta$ is the reported value. For visual simplicity, horizontal load steps are reported with an A and vertical load steps with a B. The material model parameters used for data generation were $\boldsymbol{\theta}^{true}=[42.51,13.63,14.35,11.19]$.
In reality, all 16 possible load paths would not be evaluated experimentally, since the goal of this work is to determine the one optimal path. Therefore, the load path was then chosen adaptively, as would be done in reality, with the ICC frame work by calculating the EIG of the candidate load steps in $\Xi$ at each decision point as the cruciform was actively guided through the load path tree. The EIG of each candidate load step is shown in Fig. 9 for each decision point. The EIG steadily decreases as the load steps progress, implying that additional data—for either load step option—has a waning influence on the calibration. An A load step was chosen each time for a total load path of AAAAA. As was confirmed with the pre-determined load paths in Table 4, this is in fact the load path with the least collective parameter uncertainty given the data, so while it was not guaranteed, in this case the myopic decision-making did result in the globally optimal load path (given the preset tree structure).
Figure 9: The EIG for each candidate load step in the load path.
Fig. 10 shows the posterior contours for the AAAAA load path (a), as well as a side-by-side comparison with contours for an ABABA load path (b) so that a visual comparison can be made between the cases with least and greatest parameter uncertainty. Contours for the ABABA load path are shown separately in Fig. 3 in G. The diagonal plots in (a) and (b) show the marginal posterior distribution for each parameter along with the 95% credible interval (CI) (dashed black lines) as well as $\boldsymbol{\theta}^{true}$ (pink dashed line). The off-diagonal plots show the joint marginal posterior contours for each parameter pair along with $\boldsymbol{\theta}^{true}$ indicated by a pink star. The parameter uncertainty is described by the posterior distribution, and the MAP parameter estimate corresponds with the value which maximizes the distribution.
The two load paths resulted in posterior distributions that exhibited different marginal and joint parameter uncertainty, as well as different parameter correlations (reported in E). This visual comparison makes it clear that even though both load paths resulted in expected values that were very similar (Table 4), the results of the inference are not the same. Importantly, load path AAAAA demonstrated lower parameter uncertainty compared to load path ABABA. The selection of an AAAAA load path over ABABA, or any other alternative, is not necessarily intuitive a priori. This result shows that the BOED algorithm was successfully able to identify the load path that produced more informative data than an alternative load path which may be selected a priori by subject-matter expertise.
[…]
[figure]style=plain,subcapbesideposition=top,font=small
[…]
Figure 10: The posterior contour for load path (a) AAAAA and (b) both AAAAA and ABABA. Marginal posterior densities are plotted on the diagonals and joint-marginal densities on the off-diagonals. $\boldsymbol{\theta}^{true}$ is indicated with a dashed pink line and star and the 95% CI with dashed black lines.
The posterior probability is visualised over the surrogate model output (load and PCA singular values) by drawing 200 parameter samples from the posterior distribution $\boldsymbol{\theta}^{draw}\sim\pi(\boldsymbol{\theta}\>|\>\boldsymbol{y},\boldsymbol{\xi})$ and evaluating the surrogate model at each sample $\tilde{g}(\boldsymbol{\theta}^{draw})$. The corresponding posterior probability of the load values at load step 5 is shown in Fig. 11 and for the PCA singular values in Fig. 3 in F for load path AAAAA. The model output at the MAP parameter estimate is shown with a dotted black line, and the 95% CI of the distribution over the load is shown with dashed black lines, indicating the interval in which the load values have a 95% probability of residing within.
The singular values obtained from $\tilde{g}(\boldsymbol{\theta}^{draw})$ were then used along with the 5 retained principal components to reconstruct the displacement field, $\text{disp}_{i}^{\text{recon}}=\text{disp}^{\text{PCA}}_{i}V^{*T}_{i}+\text{disp}^{\text{mean}}_{i},\,i\in[X,Y]$, so that there were 200 reconstructed instances of the displacement field. Fig. 12 shows the observed displacement field generated with $\boldsymbol{\theta}^{true}$ at the final load step for the AAAAA load path along with two line scans, across which the 95% CI was computed from the 200 reconstructed fields and plotted in Fig. 13. Given the relative scale of the displacements and the low uncertainty, the 95% CI interval appears as a single line. Overall, the high confidence in the parameter values translates to a high confidence in the resulting model output.
Figure 11: Posterior probability over the load for the AAAAA load path for both the $X$ (left column) and $Y$ (right column) directional components. The load at the MAP parameter estimate is shown with a dotted black line, and the 95% CI of the distribution over the load is shown with dashed black lines, indicating the interval in which the load values have a 95% probability of residing within.
[…]
Figure 12: The observed displacement field generated with $\boldsymbol{\theta}^{true}$ and added noise for the (a) X direction and (b) Y direction of the AAAAA load path at the final load step reported in microns. Lines (1) and (2) indicate the space over which the posterior probability of the displacement field is plotted in Fig. 13.
[…]
[]Refer to caption
[…]
Figure 13: Posterior probability over the nodal displacements for the AAAAA load path across the two lines indicated in Fig. 12. The $X$ (left column) and $Y$ (right column) directional components are shown along with Line 1 (top row) and Line 2 (bottom row). The 95% CI of the distribution over the displacement is shown with dashed black lines. Given the relative scale of the displacements and the low uncertainty, the 95% CI interval appears as a single line. The x-axis is the normalized location across each line scan.
Results for the ABABA load path calibration, which yielded the highest parameter uncertainty of the pre-determined load paths—Table 4—can be found in G. The observed data (Figs.1 & 2), along with the posterior uncertainty over the load (Fig. 4), PCA singular values (Fig. 5), and displacement field line scans (Figs. 6 & 7) are included. Similar results to the AAAAA case were observed: The posterior probability over the parameters translated to high confidence in the model output and recovered displacement field.
\@fb@secFB
4.3 Model Validation
Once the calibration is complete, the results are checked with external validation by using the model to make predictions on data not seen in calibration and comparing them to observed data. If the model obtained through the calibration fits the data well, then data generated under the model should look similar to new observed data. The new data could be taken from a new load path or specimen geometry, or it could simply be data from the same load path and geometry that could have been observed but was not (i.e. if the same experiment were to be performed again tomorrow).
Data for a new load path was predicted by drawing 200 samples from the joint posterior distribution $\pi(\boldsymbol{y}^{pred},\boldsymbol{\theta}\>|\>\boldsymbol{y})$. This is known as the posterior predictive distribution, which is the distribution of the predicted data if the calibrated model is true,
[…]
Although any of the 16 possible load paths starting with and A load step could have been chosen, an ABBBA load path was used for validation here. By comparing observed data to the posterior predictive distribution, the fit of the posterior distribution can be evaluated.
Figure 14 shows the posterior predictive distribution over the load for an ABBBA load path for the AAAAA-calibrated model. The distribution over the load is shown for each load step and for each directional component, $X$ and $Y$. Also shown is the observed data (solid red line) generated with $\boldsymbol{\theta}^{true}$ and added noise for the new load path, the predicted load at the MAP parameter estimate (dotted black line), and the 95% CI (dashed black lines). For each load step and for both directional components, the observed data is well within the 95% CI and has high probability under the assumed model.
Figure 14: The posterior predictive distribution for an ABBBA load path is shown over the load for the $X$ (left column) and $Y$ (right column) directional components and at each load step. The observed data is shown with a solid red line, the load at the MAP parameter estimate is shown with a dotted black line, and the 95% CI is shown with dashed black lines.
Visualizing the posterior predictive distribution over the displacement field is more challenging; two complementary visualizations are utilized here. Although it is challenging to visualize the uncertainty over the entire field, 3 of the 200 samples from the posterior predictive distribution were used to reconstruct corresponding displacement fields and are visualized in Fig. 15 for the X directional component. The observed data obtained at the end of each of the 5 load steps of the ABBBA loadpath is shown along with 3 samples from the predictive distribution. This allows for a comparison to be made between the displacement field that is observed and that which is expected to be observed under the posterior distribution that was obtained. The multiple predictive samples serve to establish repeatability, or that the patterns seen in the displacement field did not merely happen by chance. There is no visible systematic discrepancy between the observed and predicted data, and it is reasonable to suggest that the observed data came from the same distribution as the predictions. Similar results were seen for the Y directional component—Fig. 4 in F.
Figure 15: Samples from the posterior predictive distribution of an AAAAA load path is shown for the $X$ directional component of the displacement field for a new ABBBA load path. The left column shows the observed data for each of the 5 load steps (LS) and the three right columns show three different predictive samples.
The 95% CI of the distribution was calculated over the field point-wise from the 200 samples. Fig. 16 shows a full-field contour plot from which the 95% CI over Line 1 was extracted and plotted in Fig. 17 for all 5 load steps (LS) to show the predictive uncertainty. Plots in the left column show the 95% CI (dashed lines) along with the observed data (red line). Since distinguishing between the lines is difficult in some plots, the plots in the right column show the difference between the 95% CI and the observed data. The solid line represents the observations (centered at 0 difference), and the difference between the observations and the CIs are shown with dashed lines. Everywhere the observations fall outside the 95% CI coincides with locations where the upper bound difference is negative or the lower bound difference is positive; these regions are colored orange. Very few points along Line 1 have observations that fall outside the 95% CI. Similar observations were made for the predictive distribution over Line 2 as is shown in Fig. 5 in F.
[…]
Figure 16: The observed displacement field for the (a) X and (b) Y directional components of the ABBBA load path at the final load step reported in microns. Lines 1 and 2 mark the space over which the posterior predictive distribution is plotted in Figs. 17 and 5.
Figure 17: The posterior predictive distribution (from an AAAAA load path) over Line 1 displacements (Fig. 16) is shown for new load path ABBBA. The two left columns show the 95% CI of the distribution as well as the observations for the $X$ and $Y$ directional components for all 5 load steps (LS). The two right columns show the difference between the 95% CI and the observations. Regions where the observations fall outside the 95% CI are colored orange, and 0 is marked with a red line. the x-axis is the normalized location along Line 1.
Similar plots were made for the posterior predictive distribution from the ABABA calibration for the new ABBBA load path in G.3 to serve as a visual comparison of the two cases. Additionally, a quantitative comparison is made between the predictive accuracy of the AAAAA vs ABABA calibrations in F.4 through the log predictive density (or log likelihood). This predictive score slightly favored the AAAAA load path overall.
In summary, the model validation showed that the model obtained through the calibration fit the data well. The posterior predictive distribution was visualized over the load at each load step as well as over the displacement field. There were no systematic discrepancies seen between the observed data and the predicted data. Additionally, the observed data looked plausible under the posterior predictive distribution; generally speaking, the observed data had high probability under the assumed model and fell within the 95% CI. The predictive accuracy of the AAAAA- and ABABA-calibrated models was compared visually through plots of the predictive density over the load and displacement field as well as quantitatively through their best-fit log predictive density of 16 new data sets not seen in calibration. The results showed that the AAAAA load path generally yielded greater predictive probability than the ABABA load path. Thus, the well-calibrated and validated model with uncertainty metrics can be used for decision-making with a quantified level of confidence. Importantly, the ICC framework achieved this efficiently and with greater credibility than would have been attained through alternative load paths.
5 Discussion
Algorithmic Decisions: Assumptions and Simplifications
In the construction of the ICC framework, certain assumptions and simplifications were made to ease the burden of algorithm development. Some of these constraints may be relaxed or adjusted in future applications; however, others are necessary in order to make ICC feasible with real-time data collection and model calibration, which is the future application space of ICC. These assumptions are bulleted and discussed below along with some suggestions of relaxing these constraints in the future where possible.
  • 1.
    Posterior Approximation
    The posterior distribution was obtained with a Laplace approximation that assumes the posterior is Gaussian. Another method for approximating the posterior distribution which does not make such a strict assumption is Markov chain Monte Carlo (MCMC), which is widely used but very expensive. While the posterior is not guaranteed to be Gaussian, the Laplace approximation is obtained in less than a minute, while MCMC could easily take hours or more (conservatively) even with a fast surrogate model. Since the posterior is calculated after each load step, it is quickly realized that completing an entire experiment consisting of 5 or more load steps would not be feasible with such a computational demand. The Laplace approximation serves to quickly provide an estimate of parameter uncertainty which is then used in the EIG calculation to decide the next load step. It is possible that MCMC could result in a different adaptively selected load path if the posterior distribution is non-Gaussian; however, this possibility is secondary to efficient and practical decision-making. At the end of the data collection, MCMC could optionally be performed offline for a final calibration if desired.
  • 2.
    Measurement Error
    The measurement noise variance for both the nodal displacements as well as the global load values was assumed to be known and i.i.d. In reality, noise from a load cell is a percent of the reading instead of constant. Additionally, while noise may be approximated to mimic what is seen in experiment, the statistical model would be more flexible if the noise variance were also inferred from the data. The incorporation of unknown and hetereoscedastic noise into the BOED framework was beyond the present scope, but may be included in future work.
  • 3.
    Binary Load Path Tree
    A cruciform specimen in a planar biaxial load frame was selected for this synthetic experimental configuration. The corresponding load path tree had two branches per node, corresponding to tensile displacement in either the horizontal or vertical arms of the cruciform, with a fixed number of load steps and a fixed displacement increment for the load path tree. This was a simple starting point; however, ICC is not restricted to this tree structure. Other, more complex, and perhaps even more appropriate, branch options may include both positive and negative displacements, combinations of displacements applied in both directions simultaneously, a variable displacement increment, etc. Additionally, other deformation modes (i.e. tension-torsion) could be considered, or even a combination of planar biaxial and torsional loading given the capabilities of the load frame.
  • 4.
    Quarter symmetry of the Cruciform Specimen
    Quarter symmetry was assumed in the circular region of interest of the cruciform to reduce the FE model size and conserve computational resources. While this is a reasonable assumption for the synthetic demonstration, a real-world lab environment may present more complex behavior.
  • 5.
    Likelihood Contributions from Multiple QoIs
    Combining several QoIs for calibration is not foreign to engineering applications; both nodal displacements—transformed through PCA to yield corresponding singular values—and global load values were used to calibrate the material model in this work. However, a disparity in the number of measurements and magnitude of the values may lead to one QoI driving the likelihood and calibration disproportionately. While the number of displacement measurements ($>1,000\to 5$ after PCA) were on the same scale as the number of load measurements—1—for each directional component, there was a stark difference in the magnitude of the different QoIs—$1\times 10^{-6}-1\times 10^{-4}$ for the PCA singular values vs. $1\times 10^{3}$ N for the load. Working with kN instead of N would still present 4-6 orders of magnitude difference in the measurements. Thus, obtaining a proper balance between multiple QoIs in the likelihood is an important consideration. Table 5 shows the various contributions to the likelihood of the AAAAA load path at $\boldsymbol{\theta}^{\text{MAP}}$, revealing that the likelihood contribution from the PCA singular values is much greater than that from the load at the specified point. Future studies will further analyze the contribution of each QoI, and explore methodologies for ensuring a proper balancing of their contributions.
    $\boldsymbol{\theta}$ $\text{disp}^{\text{PCA}}_{X}$ $\text{disp}^{\text{PCA}}_{Y}$ $\text{load}_{X}$ $\text{load}_{Y}$ Log Likelihood
[…]
    Table 5: Likelihood contributions from the multiple QoIs.
  • 6.
    Myopic decision-making
    The load step selection was performed with a greedy (a.k.a myopic) decision-making protocol, where each load step was chosen based on the immediate expected gain in information—looking only one step ahead. The alternative is to make a globally optimal decision by calculating the EIG for the entire load path; however, this increases the cost of calculating the EIG exponentially. For example, at load step $t=2$, instead of calculating the EIG for 2 candidate load steps (Axis A vs. Axis B) in a mypocially-driven decision, the EIG would need to be calculated for the $n^{2}$ possible load paths which follow from the $n$ remaining number of load steps in a global decision. So, while the myopic algorithm is not guaranteed to result in the globally optimal load path, the resulting time savings make ICC possible in quasi-real-time and may yet result in the globally optimal decision (as in this example).
  • 7.
    Surrogate replacement of the FE model
    A fast Gaussian process (GP) surrogate replacement of the FE model was used for calibration and load step selection. A nice feature of GPs is that uncertainty surrounding the surrogate model can be incorporated into the inference problem. In this case, a simplified approach was taken, and the mean of the surrogate model was used without taking into account uncertainty of the surrogate model. The exclusion of the surrogate uncertainty was a simplification adopted for this initial demonstration, but future work may be expanded to include it. Additionally, because the surrogate model is trained on a grid of parameter values, the inference is restricted to this training space with truncated normal prior distributions. The parameter bounds used for training were chosen to be very conservative. However, if the support of the target posterior distribution falls outside the parameter bounds, it is likely the case that the chosen model-form is not appropriate for the specific application and should be revisited.
Ending Criteria
The ending criteria for the ICC feedback loop was to exit after a pre-determined number of load steps with the aim to keep the material sample in the plastic regime of deformation before failure. However, there are other ending criteria that could be employed. For one, the ending criteria could be the difference in the EIG between two consecutive load steps—if the EIG does not change by some threshold amount, then there is very little information expected to be gained from collecting additional data. Another approach could be to determine a threshold level of allowable uncertainty for a given application. These ending criteria are all different ways to define what is “good enough” in model calibration, which is highly application-dependent. Whatever is decided, the value of performing ICC remains the same, which is to efficiently perform model calibration for high-fidelity models with credibility that is both improved from traditional approaches and quantified.
Calibration of Multiple Models
In this phase of algorithm development, the ICC framework was constructed and implemented for the calibration of a single material model. However, the intent is to eventually use this framework for the calibration of multiple models as there is generally a level of uncertainty regarding the best model form for a given application. The Library of Advanced Materials for Engineering (LAMÉ) [34], part of the Sierra/SolidMechanics [35] finite element software, offers a modular framework for model development from which various model forms of differing degrees of fidelity may be explored. With the consideration of multiple models, there is the additional factor of determining how the step selection will proceed—this is the topic of future work.
Experimental Considerations
This work presented a demonstration of the ICC framework in a tightly-controlled synthetic environment. The synthetic demonstration was constructed to mimic the planned experiment as closely as possible, which is to deform an aluminum cruciform specimen in a planar biaxial load frame and perform calibration with full-field DIC data. A first demonstration of the ICC framework has been performed experimentally, and a primary challenge moving to the real-world laboratory conditions is machine control. The surrogate models were constructed with some assumed boundary conditions, and reliable use of these models requires that the realized experimental boundary conditions match the ones used for surrogate training as closely as possible. With some adjustments to the ICC framework [49], the strict boundary conditions were closely met and ICC was completed successfully. However, future demonstrations may require that surrogates be trained over a range of boundary conditions in order to relax the requirements on machine control. Additionally, in this synthetic demonstration, the calibration and load step selection was performed in the absence of model-form error since the same model was used to both generate the data and in calibration. In a real-world application, no model can perfectly capture the the true behavior of the material. If there is an observed systematic discrepancy between the data and the model, model-form error will need to be taken into account in order to obtain credible results in the calibration.
6 Conclusions
Computational simulation is relied upon for high-consequence engineering decisions; in solid mechanics simulations, a foundational element is a credible constitutive (material) model which can compute the stresses in a material under a specified loading condition. The usefulness of these simulations for decision-making not only relies on the credibility of the calibrated material-dependent parameters—and inherently the appropriateness of the data collected through material characterization—but also on the ability to efficiently achieve the credible calibration for the material model. Both of these considerations support the credible use of computation simulations at early stages in design cycles.
Historically, the material characterization and model calibration workflow has been sequential, inflexible, and potentially incomplete. First, relevant phenomenological features are assumed a priori based on subject matter expertise. Next, a suite of characterization experiments are designed and performed based on those assumptions. Most often, the experiments target specific stress states (i.e. tension, shear, etc.) and collect only global force-displacement data for calibration. Then, typically the material model is calibrated using a deterministic process that provides a best-fit parameter estimate with no uncertainty metrics. Finally, a distinct validation effort may—or may not—be performed. Only at the end of this process is the accuracy of the calibration ascertained. If the calibration is inadequate, then the entire process must be repeated, and it can take months or even years before a satisfactory calibration is obtained.
This work brings together several recent advancements made in material characterization and model calibration into an improved workflow—Interlaced Characterization and Calibration (ICC)—that improves not only the process of obtaining a reliable material model calibration, but also the quality of the calibration itself. The ICC framework achieves these improvements by 1) replacing uniaxial test specimens with a cruciform specimen that reveals complex stress states, 2) replacing global load-displacement curves with full-field displacement data, 3) replacing deterministic calibrations and best-fit parameter estimates with Bayesian inference, which produces a distribution quantifying parameter uncertainty, 4) adaptively selecting the load steps—within the confines of a load path tree—as the cruciform is deformed through Bayesian optimal experimental design (BOED) instead of relying on subject matter expertise to select a load path a priori, and 5) performing steps 1-4 in a quasi-real-time feedback loop. The result is that informative experimental data is collected in a highly-efficient manner to yield a credible material model that can be used for early decision-making.
A synthetic demonstration of the ICC framework was presented which considered the deformation of cruciform specimen in a biaxial load frame to calibrate an elastoplastic constitutive model. In the exemplar problem, calibration was performed with a surrogate replacement of the expensive FE model, and full-field displacement data and global load data were synthetically generated for calibration. The displacement field was dimensionally reduced via principal component analysis (PCA) to improve numerical stability and aid in the construction of surrogate models, and inference was performed with the resulting PCA singular values. A non-alternating AAAAA load path (pulling on the same axis for all five load steps) was preferred by the algorithm and resulted in a calibration that had the least parameter uncertainty compared to all other load paths—a non-intuitive result. This outcome reinforced the utility of BOED algorithms to identify load paths that collect more informative data than load paths selected a priori by subject matter experts.
The tools developed for this synthetic demonstration make a significant step in moving towards a real-time demonstration of the ICC framework in the lab. By using this approach, high-fidelity material models can be calibrated with a great increase in efficiency over current methods. Additionally, the proposed approach improves the quality of the calibrations with guided data collection and quantified uncertainty. These results improve the state-of-the-art in standard calibration procedures.
7 Acknowledgements
This work was supported by the Laboratory Directed Research and Development program at Sandia National Laboratories, a multimission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International Inc., for the U.S. Department of Energy’s National Nuclear Security Administration under contract DE-NA0003525.
This paper describes objective technical results and analysis. Any subjective views or opinions that might be expressed in the paper do not necessarily represent the views of the U.S. Department of Energy or the United States Government.
8 CRediT Authorship Contribution Statement
D.E. Ricciardi: Data curation, Formal analysis, Investigation, Methodology, Software, Validation, Visualization, Writing – original draft, Writing – review & editing. D.T. Seidl: Data curation, Methodology, Resources, Supervision, Validation, Writing – review & editing. B.T. Lester: Resources, Supervision, Writing – review & editing. A.R. Jones: Resources, Writing – review & editing. E.M.C. Jones: Conceptualization, Funding acquisition, Project administration, Supervision, Writing – review & editing.
Appendix A Cruciform Specimen
The dimensions of the cruciform geometry simulated in this work are detailed in Fig. 1.
Figure 1: Abbreviated drawing of the cruciform specimen geometry. Dimensions are in inches.
Appendix B Retained Principal Components of the Nodal Displacements
The retained principal components at Node 1 in the load path tree (Fig. 2) are shown for both directional components X and Y in Figs. 1 and 2, respectively.
[…]
Figure 1: The 5 retained principal components for the $X$ directional component of displacement after one load step along Axis A - i.e. Node 1 in the load path tree in Fig. 2
[…]
Figure 2: The 5 retained principal components for the $Y$ directional component of displacement after one load step along Axis A - i.e. Node 1 in the load path tree in Fig. 2.
Appendix C EIG Approximation
\@fb@secFB
C.1 Challenges of Calculating the EIG
In each step of the ICC framework, a single load step is taken, data is collected for calibration, Bayesian inference is performed to obtain an approximation to the posterior distribution describing parameter uncertainty, and then the next load step is selected by calculating the expected information gain (EIG). The EIG is approximated via a nested Monte Carlo estimator for each candidate design $\xi$,
$\widehat{EIG}_{MC}(\xi)=\frac{1}{N}\sum_{i=1}^{N}\log\frac{f\left(\boldsymbol{y}_{i}\>|\>\boldsymbol{\theta}_{i,0},\xi\right)}{\frac{1}{M}\sum_{j=1}^{M}f\left(\boldsymbol{y}_{i}\>|\>\boldsymbol{\theta}_{i,j},\xi\right)},\quad\boldsymbol{\theta}_{i,j}\sim\pi(\boldsymbol{\theta}),\quad\boldsymbol{y}_{i}\sim f(\boldsymbol{y}\>|\>\boldsymbol{\theta}_{i,0},\xi),$ (16)
where the denominator is a Monte Carlo approximation of the evidence,
[…]
The candidate design which yields the maximum EIG is expected to result in the most informative data for reducing parameter uncertainty.
Numerical underflow may occur when calculating the EIG as a result of one or more factors: 1) If the prior distribution is not sufficiently centered on the posterior of $\boldsymbol{\theta}_{i}$, this may lead to calculations for the evidence term (the inner loop of the EIG MC estimator) that are below machine precision [64], and 2) in the presence of high-dimensional data where the covariance-scaled squared difference of the measurements and model output (the exponential term in Eqn. (13)) are less than 1, the likelihood calculations (used to calculate the evidence) may be below machine precision. Both sources of underflow were seen in this work. The former was handled through importance sampling (details in C.2). The latter case is observed if the displacement field is used for calibration without dimension reduction; this was successfully resolved with the PCA dimension reduction process discussed in Sec. 3.2. In combination, these two approaches were successful in resolving issues with underflow.
\@fb@secFB
C.2 Importance Sampling
In the synthetic cruciform example, underflow was observed at the first adaptively selected load step (during the EIG calculation for load step two) when the sampling distribution in the EIG ($\pi(\boldsymbol{\theta})$) was not yet sufficiently centered on the posterior. As more data became available at later load steps, underflow was not observed. Furthermore, the occurrence of underflow was rare. Calculating the EIG with $M=1\times 10^{3}$ and $N=1\times 10^{4}$ resulted in three outer loop samples (.03% of the N outer loop samples) experiencing underflow for the EIG approxiamiton of an A load step, and 0 occurrences of underflow for the EIG approximation of a B load step. For a greater computational cost, increasing $M$ may aid in reducing the occurrence of underflow as more inner loop samples provide a better approximation of the evidence.
To overcome underflow when it occurred, importance sampling was used, whereby a sampling distribution alternative to the prior was chosen which contained more mass over the posterior of $\boldsymbol{\theta}_{i}$. Selection of a proper distribution to sample from is not a trivial task. Several different approaches have been proposed to calculate a sampling distribution for the EIG inner loop approximation of the evidence, two of which were studied for this work: 1) The Laplace-based importance sampling approach proposed by Beck et al. [64], and 2) the self-normalizing importance sampling approach proposed by Feng [65]. In addition to these, the analytical approach presented by Long et al. [66] was explored, which uses the Laplace approximation to estimate the evidence term of the inner loop.
It was found that the self-normalized distribution [65] was also susceptible to underflow for this problem, so this approach was not adopted. Using the Laplace-based biasing distribution [64] for all $N$ outer loop iterations would be prohibitively expensive (a single iteration of the outer loop requires 8 seconds as opposed to 8 milliseconds when sampling from the prior). Likewise, an analytical calculation of the inner loop [66] was computationally expensive. Thus, in this work, a hybrid sampling approach was used: The prior distribution was sampled from by default, but if underflow occurred, the Laplace-based importance sampling approach [64] was reverted to in order to overcome underflow for that iteration of the outer loop. Since underflow occurred for such a dearth of outer loop iterations, the hybrid approach was feasible and successfully overcame the issue with underflow with an unremarkable increase in computational time.
Appendix D Linear Transformation of a Random Variable
For some random variable $Q$ that has expected value $\mathbb{E}[Q]$ and variance $\text{Var}[Q]$, the transformation of $Q$ via $AQ$, where $A$ is a constant matrix $A\in\mathbb{R}^{i,j}$ results in  [67]
$\mathbb{E}[AQ]=A\mathbb{E}[Q],$ (18)
$\text{Var}[AQ]=A\text{Var}[Q]A^{T}.$ (19)
Nodal displacement data along with global load values were used to calibrate the material model parameters through Bayesian inference. The data $\boldsymbol{y}=[\text{disp}_{i},\text{load}_{i}]\,i\in[X,Y]$ are considered as random variables in this context. Thus, when the displacement data is dimensionally reduced—through PCA via the process discussed in Sec. 3.2—a linear transformation takes places such that $\text{disp}_{i}^{\text{PCA}}=\text{disp}_{i}V_{i}^{*}$ for $i\in[X,Y]$, where $V_{i}^{*}$ is the PCA basis of the retained principal components for the given directional component. After the transformation, $\boldsymbol{y}=[\text{disp}_{i}^{\text{PCA}},\text{load}_{i}],\,i\in[X,Y]$.
\@fb@secFB
D.1 Likelihood Construction
The likelihood for the data is modeled as a product of the likelihood of each of the 4 data sets contained in $\boldsymbol{y}$ (load and displacement for $X$ and $Y$ directional components) under the assumption of independence,
$f\left(\boldsymbol{y}\>|\>\boldsymbol{\theta},\boldsymbol{\xi}\right)=\prod_{h}\prod_{i}f\left(\boldsymbol{y}_{h,i}\>|\>\boldsymbol{\theta},\boldsymbol{\xi}\right),\quad h\in[\text{disp}^{\text{PCA}},\text{load}],\quad i\in[X,Y].$ (20)
Each data set is modeled with a multivariate normal distribution with i.i.d. Gaussian error. The multivariate normal distribution for each data set is characterized by its mean $\mathbb{E}[\boldsymbol{y}_{h,i}]=\tilde{g}(\boldsymbol{\theta})_{h,i}$ and variance $\text{Var}[\boldsymbol{y}_{h,i}]=\Psi_{h,i}$ such that $\boldsymbol{y}_{h,i}\sim\mathcal{N}\left(\mathbb{E}[\boldsymbol{y}_{h,i}],\text{Var}[\boldsymbol{y}_{h,i}]\right)=\mathcal{N}\left(\tilde{g}(\boldsymbol{\theta})_{h,i},\Psi_{h,i}\right)$. Thus, the likelihood for each independent data set is,
$f\left(\boldsymbol{y}_{h,i}\>|\>\boldsymbol{\theta},\boldsymbol{\xi}\right)$ $=\mathcal{N}_{N_{h,i}}\left(\boldsymbol{y}_{h,i};\tilde{g}(\boldsymbol{\theta})_{h,i},\Psi_{h,i}\right),$ (21)
$h\in[\text{disp}^{\text{PCA}},\text{load}],\,i\in[X,Y].$
The PCA-transformed displacement field is obtained via a linear transformation, $\text{disp}^{PCA}_{i}=\text{disp}_{i}V_{i}^{*}$, and is distributed as $\text{disp}^{PCA}_{i}\sim\mathcal{N}_{N_{\text{disp}^{PCA},i}}\left(\tilde{g}(\boldsymbol{\theta})_{\text{disp}^{\text{PCA}},i},\,V^{*T}_{i}\text{Var}[\text{disp}_{i}]V_{i}^{*}\right)$ for $i\in[X,Y]$. Recall, the surrogate models $\tilde{g}(\boldsymbol{\theta})$ take a parameter vector as input an return PCA singular values for the displacement field and load values. The noise variance for the displacement is defined in the field space, so it must be transformed according to Eqn. (19) in order to account for the PCA transformation of the displacement field.
Appendix E Parameter Correlation
See Tables 1 and 2 for posterior correlation of the AAAAA and ABABA load paths, respectively.
$\sigma_{y}$ $A$ $n$ $a$
[…]
$A$ 0.33 1.
$n$0.55 -0.64 1.
$a$0.19 0.56 0.19 1.
Table 1: Element-wise posterior correlation for load path AAAAA.
$\sigma_{y}$ $A$ $n$ $a$
[…]
$A$ 0.50 1.
$n$0.63 -0.97 1.
$a$0.17 0.14 -0.02 1.
Table 2: Element-wise posterior correlation for load path ABABA.
Appendix F Additional Figures for Load Path AAAAA Calibration
Additional figures for the load path AAAAA calibration are contained here. These include the synthetic displacement data in F.1, the posterior probability over the PCA singular values in F.2, the posterior predictive distribution over the displacement field for the ABBBA load path in F.3 and a table reporting the log predictive density for various load paths in F.4.
\@fb@secFB
F.1 Synthetic Nodal Displacement Data
See Figs. 1 and 2 for the synthetic nodal displacement data used for calibration with an AAAAA load path.
[…]
Figure 1: Synthetic DIC data for AAAAA load path. The X component of displacement is shown in microns for load steps 1-5 in plots (a)-(e), respectively.
[…]
Figure 2: Synthetic DIC data for AAAAA load path. The Y component of displacement is shown in microns for load steps 1-5 in plots (a)-(e), respectively.
\@fb@secFB
F.2 Posterior probability of PCA Singular Values
See Fig. 3 for a visualization of the posterior probability over the PCA singular values for the AAAAA load path.
Figure 3: Posterior probability over the 5 retained PCA singular values for the AAAAA load path after load step 5 for both the $X$ (top row) and $Y$ (bottom row) directional components. The singular values at the MAP parameter estimate is shown with a dotted black line, and the 95% CI of the distributions are shown with dashed black lines, indicating the interval in which the singular values have a 95% probability of residing within.
\@fb@secFB
F.3 Validation with ABBBA Load Path
Validation of the model calibrated with an AAAAA load path was performed by making predictions about a new load path ABBBA. Figs. 15 (X directional component) and 4 (Y directional component) show samples from the posterior predictive distribution Eqn. (15) over the displacement field. The left-most column shows the observed ABBBA displacement data and the next three columns show three different draws from the posterior predictive distribution.
In Fig. 5, the distribution of the field is shown over Line 2 from Fig. 16 and is represented through the 95% CI (dashed lines). The distribution is shown for all 5 load steps (LS). The left plots show the 95% CI along with the observed data, and the right plots show the difference between the 95% CI and the observed data. The solid line represents the observations (centered at 0 difference) and the difference between the observations and the CIs are shown with dashed lines. Everywhere the observations fall within the 95% CI coincides with locations where the upper bound difference is positive and the lower bound difference is negative; these regions are colored orange.
Figure 4: Samples from the posterior predictive distribution of an AAAAA load path is shown for the $Y$ directional component of the displacement field for a new ABBBA load path. The left column shows the observed data for each of the 5 load steps (LS) and the three right columns show three different predictive samples.
Figure 5: The posterior predictive distribution (from an AAAAA load path) over Line 2 displacements (Fig. 16) is shown for new load path ABBBA. The two left columns show the 95% CI of the distribution as well as the observations for the $X$ and $Y$ directional components for all 5 load steps (LS). The two right columns show the difference between the 95% CI and the observations. Regions where the observations fall outside the 95% CI are colored orange, and 0 is marked with a red line. The x-axis is the normalized location along Line 2.
\@fb@secFB
F.4 Log Predictive Density
An additional comparison can be made between the model calibration with an AAAAA load path vs an ABABA load path by comparing their predictive accuracy given a point estimate of the fitted model, $\hat{\boldsymbol{\theta}}$—namely, the MAP parameter estimate which yields the best-fit log predictive density (a.k.a. log-likelihood) [68], $\log\pi(\boldsymbol{y}^{pred}\>|\>\boldsymbol{\theta}^{\text{MAP}})$. This predictive score summarizes the fit of the model to the data and is proportional to the mean squared error in this case since a normal model and constant variance are assumed. Table 1 reports the log predictive density of the AAAAA- and ABABA-calibrated models for 16 new data sets that were not seen in calibration for the indicated load paths. For all but a few cases, the AAAAA-calibrated model yielded best-fit predictive densities that were greater than the ABABA-calibrated model.
Prediction Path Calibration Path Best-fit Log Predictive Density
$\boldsymbol{(1\times 10^{3})}$
[…]
ABABA 113,268
AAAAB AAAAA 113,907
ABABA 113,882
AAABA AAAAA 114,291
ABABA 114,257
AAABB AAAAA 114,115
ABABA 114,116
AABAA AAAAA 114,206
ABABA 114,184
AABAB AAAAA 114,346
ABABA 114,340
AABBA AAAAA 114,476
ABABA 114,463
AABBB AAAAA 114,169
ABABA 114,116
ABAAA AAAAA 113,854
ABABA 113,752
ABAAB AAAAA 114,331
ABABA 114,344
ABABA AAAAA 114,515
ABABA 114,648
ABABB AAAAA 114,293
ABABA 114,274
ABBAA AAAAA 114,518
ABABA 114,492
ABBAB AAAAA 114,173
ABABA 114,168
ABBBA AAAAA 114,153
ABABA 114,150
ABBBB AAAAA 113,997
ABABA 113,858
Table 1: Best-fit log predictive density.
Appendix G Additional Figures for Load Path ABABA Calibration
Additional figures for the load path ABABA calibration are contained here. These include the synthetic displacement data in G.1, the posterior probability in G.2 and the posterior predictive probability over the new load path ABBBA in G.3.
\@fb@secFB
G.1 Synthetic Nodal Displacement Data
See Figs. 1 and 2 for the synthetic nodal displacement data used for calibration with an ABABA load path.
[…]
Figure 1: Synthetic DIC data for ABABA load path. The X component of displacement is shown in microns for load steps 1-5 in plots (a)-(e), respectively.
[…]
Figure 2: Synthetic DIC data for ABABA load path. The Y component of displacement is shown in microns for load steps 1-5 in plots (a)-(e), respectively.
\@fb@secFB
G.2 Posterior Probability
The posterior probability contours for the ABABA load path can be found in Fig. 3 as well the posterior probability over the load in Fig. 4, PCA singular values in Fig. 5, and the displacement field in Figs. 6 & 7.
Figure 3: The posterior contour for load path ABABA. Marginal posterior densities are plotted on the diagonals and joint-marginal densities on the off-diagonals. $\boldsymbol{\theta}^{true}$ is indicated with a dashed pink line and star and the 95% CI with dashed black lines.
Figure 4: Posterior probability over the load for the ABABA load path for both the $X$ (left column) and $Y$ (right column) directional components. The load at the MAP parameter estimate is shown with a dotted black line, and the 95% CI of the distribution over the load is shown with dashed black lines, indicating the interval in which the load values have a 95% probability of residing within.
Figure 5: Posterior probability over the 5 retained PCA singular values for the ABABA load path after load step 5 for both the $X$ (top row) and $Y$ (bottom row) directional components. The singular values at the MAP parameter estimate is shown with a dotted black line, and the 95% CI of the distributions are shown with dashed black lines, indicating the interval in which the singular values have a 95% probability of residing within.
[…]
Figure 6: The observed displacement field generated with $\boldsymbol{\theta}^{true}$ and added noise for the (a) X directional component and (b) Y directional component for the ABABA load path at the final load step reported in microns. Lines (1) and (2) indicate the space over which the posterior probability of the displacement field is plotted in Fig. 7.
[…]
[]Refer to caption
[…]
Figure 7: Posterior probability over the nodal displacements for the ABABA load path across the two lines indicated in Fig. 6. The $X$ (left column) and $Y$ (right column) directional components are shown along with Line 1 (top row) and Line 2 (bottom row). The 95% CI of the distribution over the load is shown with dashed black lines. Given the relative scale of the displacements and the low uncertainty, the 95% CI interval appears as a single line. The x-axis is the normalized location across each line scan.
\@fb@secFB
G.3 Posterior Predictive Probability
Validation of the model calibrated with an ABABA load path was performed by making predictions about a new load path ABBBA. The posterior predictive distribution—Eqn. (15)—is shown over the load in Fig. 8 and over the displacement field in Figs. 9-12.
Figure 8: The posterior predictive distribution of the ABABA load path is shown for the new ABBBA load path over the global load values for the $X$ (left column) and $Y$ (right column) directional components at each load step. The observed data is shown with a solid red line, the load at the MAP parameter estimate is shown with a dotted black line, and the 95% highest density CI is shown with dashed black lines.
Figure 9: Samples from the posterior predictive distribution of the ABABA load path is shown for the $X$ directional component of the displacement field for a new ABBBA load path. The left column shows the observed data for each of the 5 load steps (LS) and the three right columns show three different predictive samples.
Figure 10: Samples from the posterior predictive distribution of the ABABA load path is shown for the $Y$ directional component of the displacement field for a new ABBBA load path. The left column shows the observed data for each of the 5 load steps (LS) and the three right columns show three different predictive samples.
Figure 11: The posterior predictive distribution from the ABABA load path is displayed over the Line 1 displacements (Fig. 16) for a new ABBBA load path. The two left columns show the 95% CI of the distribution as well as the observations for the $X$ and $Y$ directional components for all 5 load steps (LS). The two right columns show the difference between the 95% CI and the observations. Regions where the observations fall outside the 95% CI are colored orange, and 0 is marked with a red line. the x-axis is the normalized location along Line 1.
Figure 12: The posterior predictive distribution from the ABABA load path is displayed over the Line 2 displacements (Fig. 16) for a new ABBBA load path. The two left columns show the 95% CI of the distribution as well as the observations for the $X$ and $Y$ directional components for all 5 load steps (LS). The two right columns show the difference between the 95% CI and the observations. Regions where the observations fall outside the 95% CI are colored orange, and 0 is marked with a red line. the x-axis is the normalized location along Line 2.